OpenStudy (anonymous):

Solve |7 – 2x| = 4 A - x = {3/2} B - x = {0} C - x = {3/2, 11/2} D - No Solution

OpenStudy (anonymous):

|7 – 2x| = 4 7-2x =4 or 7-2x=-4 -2x=-3 or -2x=-11 x=3/2 or x=11/2